OpenStudy (anonymous):

Write the equation in slope-intercept form. y+3=3(x-1) y-5=3/4(x-5)

OpenStudy (anonymous):

y+3=3(x-1) distribute the 3 to the x and -1 y+3=3x-3 -3 -3 y=3x-6

OpenStudy (anonymous):

Do the same for the 2nd one